We Are Given That nPr=3024 and nCr=126.What Will be the Value Of "r"?
Answer by stanbon(75887)

We Are Given That nPr=3024 and nCr=126.What Will be the Value Of "r"?
----------------------
nPr = n!/(n-r)!
nCr = n!/[(n-r)1*r!]
--------
So, nPr/nCr = r!
-----
r! = 3024/126 = 24
----
r! = 1*2*3*4 24
---
r = 4
--------------
Cheers,
Stan H.
---------------